OpenStudy (legomyego180):

When finding the area between two curves, to find the bounds can you set the equations equal? The reason I ask is because it seems counter-intuitive but I remember my professor saying you could?

zepdrix (zepdrix):

Yes. Intersection is where the functions have the same (x,y) coordinate. So if you're given y=x^2+4 and y=2x+5 for example... they intersect when they have the same y value, x^2+4 = y = 2x+5 x^2+4 = 2x+5 From there, you do the algebra to figure out what x coordinate they share.

OpenStudy (samigupta8):

Obviously , we can find the bounds only and only when we equate the equations of 2 curves. The rest is well explained by zep!!

OpenStudy (legomyego180):

Thank you to you both, sorry I didnt respond back sooner, I fell asleep last night. A follow up question. I have the problem of finding the area between \[f(x)=\left\{ \left(\begin{matrix}5\sqrt{x}, 0\le x \le1 \\ 6-x^2, 1< x \le2\end{matrix}\right) \right\}\] (sorry if this looks weird, it was my best attempt at making a piecewise function) Would I set \[5\sqrt{x}=6-x^2\] to find my bounds here? Or are they already given to me in the domain of the function?

OpenStudy (legomyego180):

@agent0smith

OpenStudy (agent0smith):

This is ONE function, not two.

OpenStudy (agent0smith):

Just find the area under the given function.

OpenStudy (legomyego180):